if x+y=k touches the parabola y=x-x² then k=?

if  x+y=k touches the parabola y=x-x² then k=?

Grade:12

2 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Nikhil, 

y=k-x=x-x2, x2-2x+k=0,x=1±√(1-k), y=x-x2, dy/dx=1-2x= -1±√(1-k)=-1 because slope of x+y=k=-1, therefore,k=1

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras

vikas askiitian expert
13 years ago

x+y=k............1

y=x-x2 ...............2

from 1 and 2

k-x=x-x2

x2 -2x+k=0 .............3

if line touches the parabola then there will be only one point of contact of line with parabola..

for this condition ,in eq 3 there should be only one value of x

so discriminant =0

   b2 =4ac

  4k=4

   k=1